OpenStudy (anonymous):

Medal!!!! I have y=10x+6 I have to switch the x y and then solve for y! Please help

OpenStudy (anonymous):

in that form you have already solved for y

OpenStudy (anonymous):

Is this what you mean by switching the x and the y? y= 10x + 6 subtract 10 from both sides y -10x = 6 subract y from both sides -10x = -y + 6 divide both sides by -10 x = 1/10y -6/10 simplify x = 1/10x -3/5

OpenStudy (anonymous):

but the op said to solve for y

OpenStudy (anonymous):

Like you said earlier, y is already solved in that equation.

OpenStudy (anonymous):

it was orginally f(x)=10x+6 . he said convert it to y=10x+6 now switch the x and y and hat do you have then solve for y

OpenStudy (anonymous):

@Catseyeglint911 @jdogw